I need help with a question that says: solve each of the following for x
*1/8=4^(6x+2)
*(1/4)^x=256
*log10x=4

Answer by stanbon(48535)

solve each of the following for x
*1/8=4^(6x+2)
2^-3 = 2^(2(6x+2))
-3 = 12x+4
12x = -7
x = -7/12
----------------------
*(1/4)^x=256
2^-2x = 2^8
-2x = 8
x = -4
--------------------------

*log10x=4
Question: Does this mean log(10x) or log(base 10)x ?
===================
Cheers,
Stan H.